OpenStudy (anonymous):

Evaluate S5 for 600 + 300 + 150 + … and select the correct answer below.

OpenStudy (anonymous):

1,162.5 581.25 37.5<--my answer 18,600

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

it should be greater than 1050 600+300+150+75+37.5=?

OpenStudy (anonymous):

I suppose so? Isn't it a geometric series?

OpenStudy (anonymous):

correct.

OpenStudy (anonymous):

So, would I be correct?

OpenStudy (anonymous):

hold on.

OpenStudy (anonymous):

600{(1/2)^5-1}/(1/2-1)

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

116.5 Am I right?

zepdrix (zepdrix):

You're adding up 5 numbers. The first of which is 600. How did they then add up to 116.5? :o That's even smaller than the first number that you're adding... You silly billy -_-

OpenStudy (anonymous):

:D! Hahaha whoops! So what should I do?

zepdrix (zepdrix):

\[\large\rm S_n=a_1\frac{1-r^n}{1-r}\]So you determined that your common ratio, \(\large\rm r\), is 1/2 since you're dividing by 2 each time. So we'll plug in all of the information, \[\large\rm S_5=600\cdot \frac{1-\left(\frac{1}{2}\right)^5}{1-\frac{1}{2}}\]

zepdrix (zepdrix):

And then let mr calculator do the rest :O understand how to punch it all in?

OpenStudy (anonymous):

Not really, but I'll try :)

OpenStudy (anonymous):

So, for the fraction part I got 1.9375.

zepdrix (zepdrix):

If you want to punch it all in at once, I would try something like this,\[\large\rm 600*((1-.5\wedge5)\div(1-.5))\]Otherwise, if you like to do it piece by piece, that's fine also :)

zepdrix (zepdrix):

(I replaced 1/2 with .5 in my calculator so it's easier to deal with).

OpenStudy (anonymous):

Ok:) Then, i got 1162.5 as an answer :)

zepdrix (zepdrix):

1.9375? That sounds correct for your fractional part. Then finally 1162.5? Yayyy good job!

OpenStudy (anonymous):

Yippee!!! If, I post another one will you answer it? And thanks so much btw!!!

zepdrix (zepdrix):

I'm hungrehhh, I need to make some foods :P but ya I'll try to come look if I get a chance. You can use the @zepdrix and it'll get my attention better than the message box.

OpenStudy (anonymous):

Ok, thanks so much:)